OpenStudy (anonymous):

Evaluate the expression Ine^2 and e^(In14)

OpenStudy (raden):

Ine^2 = 2lne = 2 * 1 = ...

OpenStudy (anteater):

Using the power rule of logs: ln e^2 = 2 ln e, Since ln e = 1, 2 ln e = 2(1) = 2 e^ln 14 --> the natural log of 14 is the power you raise e to, in order to get 14. So, if you raise e to the power that you would raise it to to get 14, you get 14. e^ln 14 = 14